# Aleph-Alpha-GermanWeb: Improving German-language LLM pre-training with model-based data curation and synthetic data generation

### v2

@@ Appendix C Experiments
We trained a 1 billion parameter transformer model with 16 layers, 2,048 hidden dimensions, 32 attention heads, and 8 key-value heads. The model uses a context length of 4096 tokens and employs modern architecture choices: rotary positional embeddings (RoPE) with a base of 500,000, SwiGLU activation functions with an MLP expansion factor of 4, and RMSNorm for layer normalization.
The model was trained for 40,000 iterations with a global batch size of 512. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e-8$, and gradient clipping at 1.0. The learning rate followed a cosine decay schedule with an initial rate of 3e-4, 400 warmup steps, 4,000 cooldown steps, and minimum learning rate of 0.0. We implemented ZeRO optimization for distributed training efficiency. The model uses BFloat16 precision and was trained using a UTF-8 tokenizer with a vocabulary size of 131,072 tokens.
We used the Pharia-1 tokenizer that was trained with the Unigram algorithm using the SentencePiece library[4].
Optim. $\beta_{1}$ $\beta_{2}$ $\epsilon$ $N_{\text{Warmup}}$ $N_{\text{Cooldown}}$ learning rate
AdamW 0.9 0.95 1e-8 1% 10% 3e-4
Table 5: Parameters of the optimizer for the 1B-Llama-model.
$N_{\text{vocab}}$ $n_{\text{Layers}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{MLP}}$ Batch Size Sequence Length
131,072 16 32 2,048 8,192 512 4,096
@@
For the second part of our experiments, we trained 8B HAT models that consist of 3 consecutive transformers. In contrast to classical tokenizer-based models, HAT splits words by a rule-based splitter. Words are processed byte-wise by the smaller outer transformers, while the bigger backbone acts as the next-word predictor on a single embedding per word. The first and third models act as encoder and decoder, that transform bytes to word embeddings and vice versa. These models are connected by linear projection layers.
@@
Optim. $\beta_{1}$ $\beta_{2}$ $\epsilon$ $N_{\text{Warmup}}$ $N_{\text{Cooldown}}$ learning rate
AdamW 0.9 0.95 1e-8 1% 10% 1e-4
Table 7: Shared parameters of the optimizer for all transformer modules assembling the 8B-HAT-model.
$N_{\text{vocab}}$ $n_{\text{Layers}}^{\text{Encoder}}$ $n_{\text{Layers}}^{\text{Decoder}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{MLP}}$ Batch Size Sequence Length
256 6 4 6 768 2,047 1,024 variable
Table 8: Hyperparameters for the Encoder/ Decoder Models of 8B-HAT. Almost all parameters are the same for the Encoder and Decoder. Only the number of layers is increased to 6 for the Encoder. As we fixate the number of words per training steps in the Backbone, the Encoder and Decoder will see a variable sequence length per trainingsbatch.
$N_{\text{vocab}}$ $n_{\text{Layers}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{MLP}}$ Batch Size Sequence Length
- 32 32 4,096 11,008 1,024 3,500
@@ Appendix D Extended results
Aleph-Alpha-GermanWeb outperforms FineWeb2 across all benchmarks. Interestingly, we find that different subsets have different strengths: our synthetic data performed particularly well on MMMLU, whereas our filtered CC and high-quality classified data performed best for HellaSwag and ARC.

### v3

@@ Appendix C Experiments
We trained a 1 billion parameter transformer model with 16 layers, 2,048 hidden dimensions, 32 attention heads, and 8 key-value heads. The model uses a context length of 4096 tokens and employs modern architecture choices: rotary positional embeddings (RoPE) with a base of 500,000, SwiGLU activation functions with an MLP expansion factor of 4, and RMSNorm for layer normalisation.
The model was trained for 40,000 iterations with a global batch size of 512. We used the AdamW optimiser with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e-8$, and gradient clipping at 1.0. The learning rate followed a cosine decay schedule with an initial rate of 3e-4, 400 warmup steps, 4,000 cooldown steps, and minimum learning rate of 0.0. We implemented ZeRO optimisation for distributed training efficiency. The model uses BFloat16 precision and was trained using a UTF-8 tokeniser with a vocabulary size of 131,072 tokens.
We used the Pharia-1 tokeniser that was trained with the Unigram algorithm using the SentencePiece library Alpha (2024).
Parameter Value
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
$\epsilon$ 1e-8
$N_{\text{Warmup}}$ 1%
$N_{\text{Cooldown}}$ 10%
learning rate 3e-4
Table 5: Parameters of the AdamW optimiser for the 1B-Llama-model.
Hyperparameter Value
$N_{\text{vocab}}$ 131,072
$n_{\text{Layers}}$ 16
$n_{\text{heads}}$ 32
$d_{\text{model}}$ 2,048
$d_{\text{MLP}}$ 8,192
Batch Size 512
Sequence Length 4,096
@@
For the second part of our experiments, we trained 8B HAT models that consist of 3 consecutive transformers. In contrast to classical tokeniser-based models, HAT splits words by a rule-based splitter. Words are processed byte-wise by the smaller outer transformers, while the bigger backbone acts as the next-word predictor on a single embedding per word. The first and third models act as encoder and decoder, that transform bytes to word embeddings and vice versa. These models are connected by linear projection layers.
@@
Parameter Value
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
$\epsilon$ 1e-8
$N_{\text{Warmup}}$ 1%
$N_{\text{Cooldown}}$ 10%
learning rate 1e-4
Table 7: Shared parameters of the AdamW optimiser for all transformer modules assembling the 8B-HAT-model.
Hyperparameter Value
$N_{\text{vocab}}$ 256
$n_{\text{Layers}}^{\text{Encoder}}$ 6
$n_{\text{Layers}}^{\text{Decoder}}$ 4
$n_{\text{heads}}$ 6
$d_{\text{model}}$ 768
$d_{\text{MLP}}$ 2,047
Batch Size 1,024
Sequence Length variable
Table 8: Hyperparameters for the Encoder/ Decoder Models of 8B-HAT. Almost all parameters are the same for the Encoder and Decoder. Only the number of layers is increased to 6 for the Encoder. As we fixate the number of words per training steps in the Backbone, the Encoder and Decoder will see a variable sequence length per training batch.
Hyperparameter Value
$N_{\text{vocab}}$ -
$n_{\text{Layers}}$ 32
$n_{\text{heads}}$ 32
$d_{\text{model}}$ 4,096
$d_{\text{MLP}}$ 11,008
Batch Size 1,024
Sequence Length 3,500
@@
C.3 HQ datasets list
Books & literature (e.g., Gutenberg); legal & government (e.g., European Patent Office (EPO) and EU Parliament); news & web text (e.g., Wikinews); encyclopedic & reference (e.g., Wikipedia); scientific & technical (e.g., peS2o); and parallel & multilingual (e.g., ParaCrawl).
@@ Appendix D Extended results
GermanWeb outperforms FineWeb2 across all benchmarks. Interestingly, we find that different subsets have different strengths: our synthetic data performed particularly well on MMMLU, whereas our filtered CC and high-quality classified data performed best for HellaSwag and ARC.